An Airplane flies a roundtrip consisting of L miles each way. The average speed on the outward leg is 160 Miles per hour, and the average speed on the return leg is 240 miles per hour. Find the average speed for the entire trip.
I think that..
R=DT
160=100(T) so T=160/100
R=DT
240=100(T) so T=240/100...so, what now?

Answer by ptaylor(1894)

HEY, YOU GAVE IT A SHOT!!!!! TRY TO FOLLOW THIS:
Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Average Speed=(Total Distance)/(Total Time)
Time required on the outward leg=L/160
Time required on the return leg=L/240
Total Distance=2L
Total Time=L/160 + L/240=3L/480 + 2L/480=5L/480=L/96, so:
Average Speed=2L/(L/96)=(2L*96)/L=192 mph--------------ans

Hope this helps----ptaylor